# VICON: Vision In-Context Operator Networks for Multi-Physics Fluid Dynamics Prediction

### v4

@@ 5 Experimental Results
We benchmark VICON against two state-of-the-art sequence-to-sequence models: DPOT [27] (122M parameters) and MPP [26] (AViT-B architecture, 116M parameters). The vanilla ICON, which would require processing over 114K tokens per frame (128×128×7), exceeds our available GPU memory and is thus computationally infeasible for direct comparison on these datasets.
@@
For conciseness, we present summarized plots and tables in the main text, and defer complete results and visualizations to appendix˜D. Ablation studies examining key design choices of VICON—including patch resolution, positional encoding, and context length—are presented in section˜D.1. The checkerboard artifacts (e.g., in fig.˜9) of the PDEBench-Comp-HighVis dataset are discussed in section˜D.2.
@@
Table 1: Summary of Rollout Relative $L^{2}$ Error (scaled by std) across different methods and datasets. The best results are highlighted in bold. For flexible step rollout, step 3 works best for the PDEArena-Incomp dataset.
Last step
[1e-2]
PDEArena-Incomp 76.77 68.03 65.27 93.52
PDEBench-Comp-LowVis 39.11 39.11 48.92 65.32
PDEBench-Comp-HighVis 61.41 61.41 2866 185.3
All average
[1e-2]
PDEArena-Incomp 56.27 48.50 41.20 55.95
PDEBench-Comp-LowVis 27.08 27.08 37.72 46.68
PDEBench-Comp-HighVis 30.06 30.06 821.9 72.37
@@ 5.1 Superior Performance on Long-term Rollout Predictions
As demonstrated in fig.˜2(a), table˜1, and fig.˜3, VICON consistently outperforms baseline methods on long-horizon predictions across all benchmarks—with the exception of DPOT on the PDEArena-Incomp dataset, where our performance is comparable. Overall, VICON achieves an average reduction in relative $L^{2}$ RMSE at the final timestep of 37.9% compared to DPOT and 44.7% compared to MPP.
@@
More detailed results are provided in table˜11 and table˜12 in appendix˜D.
@@ 5.2 Flexible Rollout Strategies
As demonstrated in fig.˜2(b) and fig.˜5, VICON can select appropriate timestep strides based on each dataset’s characteristics. While PDEBench-Comp-LowVis and PDEBench-Comp-HighVis perform best with single-step rollout, PDEArena-Incomp achieves optimal results with a stride of 3 using our flexible-step rollout strategy.
@@ 5.3 Importance of In-Context Learning
Model Stride 1 Error Stride 2 Error
@@
ViT (Last step) 73.94 95.03
ViT (All average) 50.17 63.15
@@ 5.4 Robustness to Imperfect Temporal Measurements
Real-world experimental measurements frequently suffer from imperfections—sampling rates may differ from training set, and frames can be missing due to device errors. Sequence-to-sequence models like MPP and DPOT require either retraining with new data or interpolation that introduces noise. In contrast, VICON elegantly handles such imperfections by forming context pairs from available frames only (fig.˜1(c)). The algorithm for generating pairs with imperfect measurements appears in section˜C.2.
We evaluate VICON against baselines on all benchmarks under two scenarios: (1) half sampling rate (dropping every other frame) and (2) random frame dropping (removing 1-3 frames per sequence). Results for the PDEBench-Comp-LowVis dataset in fig.˜2(c) (with complete results in fig.˜8) show that VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% in baselines—demonstrating remarkable robustness to measurement imperfections.
@@ 5.5 Turbulence Kinetic Energy Analysis
Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of [49]), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON, DPOT, and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming MPP’s error of $0.049$, while achieving performance comparable to DPOT’s error of $0.012$. Visualizations of TKE errors are presented in fig.˜6 in appendix˜D.
@@ 5.6 Benefits of Multi-Physics Joint Training
As shown in fig.˜7, joint training significantly outperforms separate training across all three datasets. We attribute this performance gain to the underlying physical similarities across these fluid dynamics problems, where exposure to diverse yet related flow patterns enhances the model’s generalization ability.
@@ 5.7 Computational Efficiency
In fig.˜2(d) and table˜14, we compare the computational resources required by VICON and baselines. Our model demonstrates superior efficiency across multiple metrics. Compared to MPP, VICON requires approximately one-third of the inference time per frame while using 75% of the total parameters. VICON also outperforms DPOT, requiring 28% fewer parameters and 28% less inference time.
@@ 6 Conclusion and Future Work
Furthermore, the current approach does not yet extend to 3D applications, as token sequence length would grow cubically, exceeding our computational budget. Potential remedies for 3D scaling include: (1) using VAE-like techniques/transformers to compress and evolve the fields in the latent space, then applying attention mechanisms to decode at any location [50, 14], and (2) employing crop-like methods to learn evolution of fixed-width windows similar to CNNs. These approaches could potentially be combined to address the cubic scaling challenge.
Since VICON is an autoregressive model, it still suffers from error accumulation of the final pair’s condition frame during rollout. Future work could explore resolving error accumulation through techniques such as injecting Gaussian noise or using push-forward methods [51].
The channel-union approach is also limited when incorporating domains beyond fluid dynamics with fundamentally different state variables. Recent advances in inter-channel attention mechanisms [52] offer promising directions for handling channel scaling issues when expanding to more diverse physics systems.
Finally, adapting VICON to handle irregular domains such as graphs or meshes would broaden its applications to areas like solid mechanics and molecular dynamics. The original ICON framework has demonstrated success across various problem types including inverse problems and steady-state tasks [25], suggesting potential for extending VICON beyond forward temporal prediction. Addressing these challenges will advance the promising paradigm of in-context learning for a broader range of physical systems.
@@ A.1 PDEArena-Incomp Dataset
The incompressible Navier-Stokes dataset comes from PDEArena [1]. The data are generated from the equation
@@
The space-time domain is $0,32]^{2}\times[18,105]$ where $dt=1.5$ and the space resolution is $128\times 128$. A scalar particle field is being transported with the fluids. The velocity fields satisfy Dirichlet boundary condition, and the scalar field satisfies Neumann boundary condition. The forcing term $\bm{f}$ is randomly sampled. The quantities of interest are the velocities and the scalar particle field.
@@ A.2 PDEBench-Comp-HighVis and PDEBench-Comp-LowVis Datasets
The PDEBench-Comp-HighVis and PDEBench-Comp-LowVis datasets come from PDEBench Compressible Navier-Stokes dataset [49]. The data are generated from the equation
@@ B.2 Training Details
We implement our method in PyTorch [53] and utilize data parallel training [54] across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in table˜5.
@@ B.3 MPP Details
For a fair comparison, we retrained MPP [26] using our training dataset using the same model configurations and optimizer hyperparameters (batch size is set to be the maximum possible on our device). We evaluate MPP using the same testing setup and metric.
@@ B.4 DPOT Details
For a fair comparison, we retrain DPOT [27] using our training dataset using the same model configurations and optimizer hyperparameters (batch size is set to be the maximum possible on our device). We evaluate DPOT using the same testing setup and metric.
@@ 30   Pdt $\leftarrow$ P[s]
Tables 8 and 9 illustrate our adaptive strategies when frames 2, 5, and 9 are missing from the initial sequence (with all other settings identical to those in section˜C.1). The tables demonstrate single-step and flexible-step (max stride: 3) approaches, respectively, highlighting how our algorithm dynamically selects appropriate example pairs to maintain prediction capability despite missing some temporal samplings.
@@ D.1 Ablation Studies
Figure 4: Ablation studies across the three datasets. Top row (a-c): Impact of patch resolutions ($8,16,32,64$) showing optimal performance at patch size 16. Middle row (d-f): Effect of different positional encoding combinations. Bottom row (g-i): Performance variation with different context lengths (6, 8, 10 pairs).
@@
We conducted ablation studies on patch resolution by varying patch sizes (8, 16, 32, 64) to find a balance between spatial granularity and computational resource constraint. While smaller patches theoretically capture finer details, they generate longer token sequences, hitting memory caps due to transformer’s quadratic complexity. For patches smaller than $16\times 16$, we had to reduce token dimensions (1024→512) and feedforward dimensions (from 2048 to 1024), while patches below $8\times 8$ remained infeasible even with mixed precision training. fig.˜4(a-c) shows that patch size 16 achieves optimal performance across all datasets—balancing sequence length and representational capacity. Performance degradation with coarser patches (32, 64) aligns with expectations, while degradation with finer patches stems from the necessary reduction in hidden dimensions, highlighting a fundamental challenge in scaling to 3D applications.
@@
We evaluated different combinations of positional encodings as shown in fig.˜4(d-f). Our architecture employs two types of encodings: patch position encodings (for spatial relationships between patches) and function encodings (differentiating input/output in different pairs) (section˜4.2).
@@
The impact of patch encodings varies across datasets: they show minimal effect on the PDEBench datasets (PDEBench-Comp-LowVis and PDEBench-Comp-HighVis) yet significantly improve performance on PDEArena-Incomp. We attribute this to the different stiffness of these systems. In compressible flows (PDEBench), spatial interactions naturally decay with distance, creating consistent and monotonic spatial correlations between patches. In contrast, Navier-Stokes equations exhibit infinite stiffness where correlations between any spatial locations are instantaneous and determined by the global velocity field, creating non-monotonic and case-dependant relationships. Explicitly encoding spatial positions therefore becomes particularly beneficial for PDEArena-Incomp, as it reduces the learning complexity for these non-local dependencies.
@@
We investigated how the number of in-context examples affects model performance, as shown in fig.˜4(g-i). Following insights from the original ICON work [25], we evaluated context lengths of 6, 8, and 10 pairs, balancing performance against computational efficiency. The results show that 10 in-context examples deliver significantly better results for PDEBench-Comp-LowVis compared to shorter contexts, while for the remaining datasets, 10 and 8 in-context examples perform similarly (both outperforming 6 examples).

### v5

@@ 5 Experimental Results
We benchmark VICON against two state-of-the-art sequence-to-sequence models: DPOT Hao et al. (2024) (122M parameters) and MPP McCabe et al. (2023) (AViT-B architecture, 116M parameters), as well as Specialist U-Net baselines Gupta and Brandstetter (2022) trained individually per dataset. Implementation details for all baselines are provided in Appendix˜B. The vanilla ICON, which would require processing over 114K tokens per frame (128×128×7), exceeds our available GPU memory and is thus computationally infeasible for direct comparison on these datasets.
@@
For conciseness, we present summarized plots and tables in the main text, and defer complete results and visualizations to Appendix˜D. Ablation studies examining key design choices of VICON—including patch resolution, positional encoding, and context length, as well as an alternative CNN-based architecture—are presented in Section˜D.1. The checkerboard artifacts (e.g., in Figure˜9) of the PDEBench-Comp-HighVis dataset are discussed in Section˜D.2.
@@
Table 1: Summary of Rollout Relative $L^{2}$ Error (scaled by std) across different methods and datasets. The best results are highlighted in bold. For flexible step rollout, step 3 works best for the PDEArena-Incomp dataset. Specialist U-Net is trained individually per dataset with instance normalization matching VICON’s preprocessing.
Rollout Relative $L^{2}$ Error Case Ours (single step) Ours (flexible step) DPOT MPP Specialist U-Net
Last step [1e-2] PDEArena-Incomp 76.77 68.03 65.27 93.52 48.8
PDEBench-Comp-LowVis 39.11 39.11 48.92 65.32 55.9
PDEBench-Comp-HighVis 61.41 61.41 2866 185.3 39.9
All average [1e-2] PDEArena-Incomp 56.27 48.50 41.20 55.95 27.9
PDEBench-Comp-LowVis 27.08 27.08 37.72 46.68 35.7
PDEBench-Comp-HighVis 30.06 30.06 821.9 72.37 19.9
@@ 5.1 Superior Performance on Long-term Rollout Predictions
As demonstrated in Figure˜2(a), Table˜1, and Figure˜3, VICON consistently outperforms baseline methods on long-horizon predictions across all benchmarks—with the exception of DPOT on the PDEArena-Incomp dataset, where our performance is comparable. Overall, VICON achieves an average reduction in relative $L^{2}$ RMSE at the final timestep of 37.9% compared to DPOT and 44.7% compared to MPP.
@@
To further contextualize VICON’s performance, we compare against Specialist U-Net baselines (U-Net-Mod64 from PDEArena Gupta and Brandstetter (2022)) trained individually per dataset. The only modification we made is adding instance normalization matching VICON’s preprocessing, which we found essential for stabilizing training on PDEBench-Comp-LowVis and PDEBench-Comp-HighVis datasets. As shown in Table˜1, while the specialist U-Net achieves strong performance on individual datasets it was optimized for, VICON remains competitive across all datasets and even outperforms the specialist on PDEBench-Comp-LowVis, demonstrating the advantage of our multi-physics approach in handling diverse physical systems without per-dataset architectural engineering or retraining.
More detailed results are provided in Table˜13 and Table˜14 in Appendix˜D.
@@ 5.2 Flexible Rollout Strategies
As demonstrated in Figure˜2(b) and Figure˜5, VICON can select appropriate timestep strides based on each dataset’s characteristics. While PDEBench-Comp-LowVis and PDEBench-Comp-HighVis perform best with single-step rollout, PDEArena-Incomp achieves optimal results with a stride of 3 using our flexible-step rollout strategy.
@@ 5.3 Importance of In-Context Learning
Model Param Size Stride 1 Error Stride 2 Error
@@
VICON (Last step) 88 M 76.77 68.03
ViT (Last step) 162 M 73.94 95.03
VICON (All average) 88 M 56.27 48.50
ViT (All average) 162 M 50.17 63.15
@@ 5.4 Robustness to Imperfect Temporal Measurements
Real-world experimental measurements frequently suffer from imperfections—sampling rates may differ from the training set, and frames can be missing due to device errors. Sequence-to-sequence models like MPP and DPOT require either retraining with new data or interpolation that introduces noise. In contrast, VICON elegantly handles such imperfections by forming context pairs from available frames only (Figure˜1(c)). The algorithm for generating pairs with imperfect measurements appears in Section˜C.2.
We evaluate VICON against baselines on all benchmarks under two scenarios: (1) half sampling rate (dropping every other frame) and (2) random frame dropping (removing 1-3 frames per sequence). Results for the PDEBench-Comp-LowVis dataset in Figure˜2(c) (with complete results in Figure˜8) show that VICON experiences only 24.41% relative performance degradation compared to 71.37%-74.49% in baselines—demonstrating remarkable robustness to measurement imperfections.
@@ 5.5 Turbulence Kinetic Energy Analysis
Within our datasets, only a subset of PDEBench-Comp-LowVis, specifically those initialized with fully developed turbulent fields (see Appendix D.5 of Takamoto et al. (2022)), is suitable for TKE analysis. After filtering these entries, we compare the mean absolute error (MAE) of the TKE between VICON, DPOT, and MPP. Our approach achieves a TKE error of $0.016$, significantly outperforming MPP’s error of $0.049$, while achieving performance comparable to DPOT’s error of $0.012$. Visualizations of TKE errors are presented in Figure˜6 in Appendix˜D.
@@ 5.6 Benefits of Multi-Physics Joint Training
As shown in Figure˜7, joint training significantly outperforms separate training across all three datasets. We attribute this performance gain to the underlying physical similarities across these fluid dynamics problems, where exposure to diverse yet related flow patterns enhances the model’s generalization ability.
5.7 Context Sampling Strategy
By default, VICON forms context pairs from initial frames within a single trajectory, which we term intra-trajectory context. This design choice is motivated by practical online deployment scenarios where only a few initial measurements are available, and system parameters (e.g., viscosity, forcing terms) may be unknown—making retrieval of historical trajectories with the same underlying operator infeasible.
As a theoretical question, we additionally investigate whether VICON can leverage cross-trajectory context, i.e., context pairs from different trajectories governed by the same operator. As detailed in Section˜D.3, VICON trained with cross-trajectory context achieves similar performance with mixed contexts during rollout (only 0.012 increase in average error). This demonstrates the model’s capability to leverage cross-trajectory information when available. Nevertheless, we argue that intra-trajectory context remains more practical as it avoids the need for maintaining trajectory databases and online operator identification.
@@ 5.8 Computational Efficiency
In Figure˜2(d) and Table˜16, we compare the computational resources required by VICON and baselines. Our model demonstrates superior efficiency across multiple metrics. Compared to MPP, VICON requires approximately one-third of the inference time per frame while using 75% of the total parameters. VICON also outperforms DPOT, requiring 28% fewer parameters and 28% less inference time.
@@ 6 Conclusion and Future Work
Furthermore, the current approach does not yet extend to 3D applications, as token sequence length would grow cubically, exceeding our computational budget. Potential remedies for 3D scaling include: (1) using VAE-like techniques/transformers to compress and evolve the fields in the latent space, then applying attention mechanisms to decode at any location Wang et al. (2024b); Li et al. (2023), and (2) employing crop-like methods to learn evolution of fixed-width windows similar to CNNs. These approaches could potentially be combined to address the cubic scaling challenge.
Since VICON is an autoregressive model, it still suffers from error accumulation of the final pair’s condition frame during rollout. Future work could explore resolving error accumulation through techniques such as injecting Gaussian noise or using push-forward methods Brandstetter et al. (2021).
The channel-union approach is also limited when incorporating domains beyond fluid dynamics with fundamentally different state variables. Recent advances in inter-channel attention mechanisms Holzschuh et al. (2025) offer promising directions for handling channel scaling issues when expanding to more diverse physics systems.
Finally, adapting VICON to handle irregular domains such as graphs or meshes would broaden its applications to areas like solid mechanics and molecular dynamics. Techniques such as geometric deep learning extensions, masked patch modeling, or point-cloud modeling are potential future directions to handle irregular geometries. The original ICON framework has demonstrated success across various problem types including inverse problems and steady-state tasks Yang et al. (2023a), suggesting potential for extending VICON beyond forward temporal prediction. Addressing these challenges will advance the promising paradigm of in-context learning for a broader range of physical systems.
@@ A.1 PDEArena-Incomp Dataset
The incompressible Navier-Stokes dataset comes from PDEArena Gupta and Brandstetter (2022). The data are generated from the equation
@@
The space-time domain is $0,32]^{2}\times[18,105]$ where $dt=1.5$ and the space resolution is $128\times 128$. A scalar particle field is being transported with the fluids. The velocity fields satisfy Dirichlet boundary conditions, and the scalar field satisfies Neumann boundary conditions. The forcing term $\bm{f}$ is randomly sampled. The quantities of interest are the velocities and the scalar particle field.
@@ A.2 PDEBench-Comp-HighVis and PDEBench-Comp-LowVis Datasets
The PDEBench-Comp-HighVis and PDEBench-Comp-LowVis datasets come from PDEBench Compressible Navier-Stokes dataset Takamoto et al. (2022). The data are generated from the equation
@@ B.2 Training Details
We implement our method in PyTorch Paszke et al. (2019) and utilize data parallel training Li et al. (2020) across two NVIDIA RTX 4090 GPUs. We employ the AdamW optimizer with a cosine learning rate schedule that includes a linear warmup phase. We apply gradient clipping with a maximum norm of 1.0 to ensure training stability. All optimization parameters are detailed in Table˜5.
@@ B.3 MPP Details
For a fair comparison, we retrain MPP McCabe et al. (2023) using our training dataset with the same model configurations and optimizer hyperparameters (batch size is set to the maximum possible on our device). We evaluate MPP using the same testing setup and metric.
@@ B.4 DPOT Details
For a fair comparison, we retrain DPOT Hao et al. (2024) using our training dataset with the same model configurations and optimizer hyperparameters (batch size is set to the maximum possible on our device). We evaluate DPOT using the same testing setup and metric.
B.5 Specialist U-Net Details
We use U-Net-Mod64, the strongest baseline from the PDEArena benchmark Gupta and Brandstetter (2022), trained individually per dataset. This architecture employs a U-Net with modified residual blocks and 64 base channels. We keep the architecture consistent across all datasets, only adjusting the input/output channel dimensions to match each dataset’s state variables. We believe this is fair because in realistic deployment, one cannot architecturally re-engineer a model for every new measurement stream encountered.
Additionally, we include instance normalization matching VICON’s preprocessing (channel-wise normalization using statistics computed from initial frames), which we found essential for stabilizing training on the PDEBench datasets (PDEBench-Comp-LowVis and PDEBench-Comp-HighVis). Without this normalization, training on compressible flows diverges due to the large magnitude variations in the pressure channel. The batch size is set to the maximum possible on our device. Each specialist model is trained on its respective dataset only, in contrast to VICON’s joint multi-physics training.
@@ 30   Pdt $\leftarrow$ P[s]
Tables 8 and 9 illustrate our adaptive strategies when frames 2, 5, and 9 are missing from the initial sequence (with all other settings identical to those in Section˜C.1). The tables demonstrate single-step and flexible-step (max stride: 3) approaches, respectively, highlighting how our algorithm dynamically selects appropriate example pairs to maintain prediction capability despite missing temporal data.
@@ D.1 Ablation Studies
Figure 4: Ablation studies across the three datasets. Top row (a-c): Impact of patch resolutions ($4,8,16,32,64$) showing optimal performance at patch size 16. Middle row (d-f): Effect of different positional encoding combinations. Bottom row (g-i): Performance variation with different context lengths (6, 8, 10 pairs).
@@
We conducted ablation studies on patch resolution by varying patch sizes (4, 8, 16, 32, 64) to find a balance between spatial granularity and computational resource constraint. While smaller patches theoretically capture finer details, they generate longer token sequences, hitting memory caps due to transformer’s quadratic complexity. For patch size 8, we reduced token dimensions (512→256) and feedforward dimensions (1024→512). For patch size 4, we applied further reductions (token dim 128, feedforward dim 256, layers 10→5) to fit within 80GB A100 memory—yet training still required 188 hours (vs 58 hours for ps=16) and inference slowed to 9.55 sec/step (vs 8.7 ms for ps=16). Patches below $4\times 4$ remained infeasible even with these aggressive reductions. Figure˜4(a-c) shows that patch size 16 achieves optimal performance across all datasets—balancing sequence length and representational capacity. Performance degradation with coarser patches (32, 64) aligns with expectations, while degradation with finer patches stems from the necessary reduction in hidden dimensions, highlighting a fundamental challenge in scaling to 3D applications.
@@
We evaluated different combinations of positional encodings as shown in Figure˜4(d-f). Our architecture employs two types of encodings: patch position encodings (for spatial relationships between patches) and function encodings (differentiating input/output in different pairs) (Section˜4.2).
@@
The impact of patch encodings varies across datasets: they show minimal effect on the PDEBench datasets (PDEBench-Comp-LowVis and PDEBench-Comp-HighVis) yet significantly improve performance on PDEArena-Incomp. We attribute this to the different stiffness of these systems. In compressible flows (PDEBench), spatial interactions naturally decay with distance, creating consistent and monotonic spatial correlations between patches. In contrast, Navier-Stokes equations exhibit infinite stiffness where correlations between any spatial locations are instantaneous and determined by the global velocity field, creating non-monotonic and case-dependent relationships. Explicitly encoding spatial positions therefore becomes particularly beneficial for PDEArena-Incomp, as it reduces the learning complexity for these non-local dependencies.
@@
We investigated how the number of in-context examples affects model performance, as shown in Figure˜4(g-i). Following insights from the original ICON work Yang et al. (2023a), we evaluated context lengths of 6, 8, and 10 pairs, balancing performance against computational efficiency. The results show that 10 in-context examples deliver significantly better results for PDEBench-Comp-LowVis compared to shorter contexts, while for the remaining datasets, 10 and 8 in-context examples perform similarly (both outperforming 6 examples).
Alternative Architecture: CNN-ICON.
We explored an alternative architecture (CNN-ICON) that uses a CNN encoder to compress entire frames into single latent vectors, followed by the ICON transformer structure. The encoder consists of a 4-level CNN with residual blocks (7×128×128 → 512-dim latent via global average pooling), paired with a symmetric CNN decoder. This results in 1 token per frame compared to VICON’s 64 tokens (8×8 patches). We matched the model sizes: 86.5M parameters (vs VICON’s 87.8M), identical transformer configuration (6 layers, 8 heads, dim=512), and same training protocol (200K steps).
Table 10: CNN-ICON vs VICON comparison (normalized RMSE, scaled by std)
Dataset Model Step 1 Step 5 Step 10 Last Avg
PDEArena-Incomp CNN-ICON 0.323 0.467 0.643 1.194 0.912
VICON 0.110 0.206 0.305 0.680 0.485
PDEBench-Comp-LowVis CNN-ICON 0.750 0.875 1.012 1.013 0.905
VICON 0.154 0.245 0.375 0.391 0.271
PDEBench-Comp-HighVis CNN-ICON 0.390 0.767 1.939 2.364 1.162
VICON 0.052 0.163 0.498 0.614 0.301
[…]